# Optimal Distributed Similarity Estimation for Unitary Channels

### v1

@@ Appendix A Preliminaries
In the following, we introduce two central tools used in this work: the Haar measure and approximate unitary designs. Both play fundamental roles in quantum information theory [50, 47], and they are essential for establishing the query complexities of our algorithms as well as the lower bounds for completing DSEU.
A.1 Haar Measure and Permutation Operators
@@
Let $\mathcal{H}$ be a Hilbert space with dimension $d$ and $\mathcal{S}_{k}$ be the symmetric group and $A\in\mathcal{B}(\mathcal{H}^{\otimes k})$. Then, we have
@@
For quantum states in $\mathcal{H}$, we define the Haar measure on quantum states as [50],
$\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}\psi^{\otimes k}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}\phi^{\otimes k}U^{\dagger\otimes k}=\frac{1}{\kappa_{k}}\Pi_{\rm sym}^{(d,k)},\quad\phi\in\mathcal{D}(\mathcal{H}),$ (20)
@@
$\mathbb{E}_{|{\psi}\rangle\sim\mathcal{A}}\psi^{\otimes k}=\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}\psi^{\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}}.$ (22)
@@
A.2 Approximate Unitary Designs
@@
Appendix B Upper Bound
B.1 Incoherent Access: Proof of Theorem 1
The algorithm solving DSEU with incoherent access is summarized in Algorithm 1. In the following, we will prove Theorem 1.
Input: number of SPAM settings $T$,
    number of measurements for each SPAM setting $m$,
    TmTm queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
1 for $t=1,\cdots,T$ do
2   Randomly generate pure states $|{\psi_{t}}\rangle$ from a 44-design state ensemble on each quantum devices.
3  Apply the unitary channels $\mathcal{U}$ and $\mathcal{V}$ on $|{\psi_{t}}\rangle$.
4  Sample a random unitary $Q_{t}$ from an unitary $4$-design ensemble.
5  Measure $m$ copies of $\mathcal{U}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{a}\rangle\!\langle{a}|Q_{t}\}_{a}$ and obtain $A=\{a_{i}\}_{i=1}^{m}$.
6  Measure $m$ copies of $\mathcal{V}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{b}\rangle\!\langle{b}|Q_{t}\}_{b}$ and obtain $B=\{b_{i}\}_{i=1}^{m}$.
7  Compute the $\tilde{g}_{t}=\sum_{i,j=1}^{m}\mathbb{1}(a_{i},b_{j})/m^{2}$, where $\mathbb{1}(a,b)=1$ if $a=b$; otherwise 0.
8  Compute
$\tilde{\omega}_{t}=\frac{(d+1)^{2}}{d}\tilde{g}_{t}-\frac{d+2}{d}.$ (28)
Algorithm 1 Distributed Similarity Estimation for Unitary Channels with Incoherent Access
$p_{Q,\psi}(a)\mathrel{\mathop{\ordinarycolon}}=\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle,\quad q_{Q,\psi}(b)\mathrel{\mathop{\ordinarycolon}}=\langle{b}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{b}\rangle,$ (29)
and their classical inner product function
Then, the expectation of $\tilde{g}$ can be represented as
$\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}$ $=\mathbb{E}_{Q,\psi}g(Q,\psi)$ (31)
$=\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (32)
$=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(\sum_{a}\mathbb{E}_{Q}Q^{\dagger\otimes 2}|{aa}\rangle\!\langle{aa}|Q^{\otimes 2}\right)\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\otimes V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right]$ (33)
$=\frac{1}{d+1}+\frac{1}{d+1}\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right],$ (34)
where we have used Lemma 7 and the property of unitary $2$-design. For the last term, we have
$\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right]$ $=\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)\left(\mathbb{E}_{\psi}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right)\right]$ (35)
$=\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)\frac{\Pi_{\rm sym}^{(d,2)}}{\kappa_{2}}\right]=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}.$ (36)
Therefore, we have
$\quad\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}=\mathbb{E}_{Q,\psi}g(Q,\psi)=\frac{1}{d+1}+\frac{1}{d}\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)},$ (37)
$\Rightarrow$ $\quad\frac{(d+1)^{2}}{d}\mathbb{E}\tilde{g}-\frac{d+2}{d}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}$ (38)
$\Rightarrow$ $\quad\mathbb{E}\tilde{\omega}_{t}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}.$ (39)
Now, we consider the variance of estimator so that we can obtain the query complexity of Algorithm 1. The following analysis is similar to [15]. With the definition of $\omega$}, the variance of $\omega$} is given by
[…]
Lemma 9.
Proof of Lemma 9.
It should be note that $\tilde{g}$ is a function of SPAM setting $\{Q,\psi\}$ and measurement results $S\mathrel{\mathop{\ordinarycolon}}=\{A,B\}$. Thus, we can write $\tilde{g}$ as $\tilde{g}(Q,\psi,S)$ and the law of total variance gives [15]
$\mathbb{V}(\tilde{g}(Q,\psi,S))=\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]\mathbb{V}_{Q,\psi}+\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right].$ (42)
We consider this two terms respectively as follows.
  1. 1.
    For the first term, with Lemma 14 in [15], we have
    As shown in Eq. (37), we have
    $\mathbb{E}_{Q,\psi}\frac{g(Q,\psi)}{m^{2}}=\frac{1}{m^{2}(d+1)}+\frac{1}{m^{2}d}\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}=\mathcal{O}\left(\frac{1}{m^{2}d}\right).$ (44)
    Additionally, with Eq. (187) in [15], we have
    $\mathbb{E}_{Q}p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)$ $=\mathbb{E}_{Q}\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle^{2}\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (45)
    $=d\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{\phi}\rangle=\mathcal{O}\left(\frac{1}{d^{2}}\right),$ (46)
    where we use the property of unitary $3$-design. Thus, we have
    Therefore, we have
  2. 2.
    For the second term, we have
    $=\mathbb{E}_{Q,\psi}\;g^{2}(Q,\psi)-\mathbb{E}_{\psi}\left(\frac{1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)]}{d+1}\right)^{2}.$ (50)
    We have
    $\mathbb{E}_{Q}\;g^{2}(Q,\psi)=\mathbb{E}_{Q}\;\left(\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\right)^{2}$ (51)
    $=$ $\;d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}+d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|\mathcal{U}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|\mathcal{V}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle,$ (52)
    $d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}=d\mathcal{O}\left(\frac{1}{d^{4}}\right)=\mathcal{O}\left(\frac{1}{d^{3}}\right),$ (53)
    $d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|\mathcal{U}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|\mathcal{V}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle=\frac{(1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)])^{2}}{d(d+1)}+\mathcal{O}\left(\frac{1}{d^{4}}\right).$ (54)
    Therefore, we have
    $\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $\leq\mathcal{O}\left(\frac{1}{d^{3}}\right)+\frac{1}{d(d+1)^{2}}\mathbb{E}_{\psi}\left(1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)]\right)^{2}=\mathcal{O}\left(\frac{1}{d^{3}}\right).$ (55)
Combining the above two terms, we have
B.2 Coherent Access: Proof of Theorem 2
The algorithm solving DSEU with coherent access is summarized in Algorithm 2. In the following, we will prove Theorem 2, which concludes the query complexity of Algorithm 2.
Input: $T$ queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
1
2Randomly generate pure states $|{\psi}\rangle^{\otimes T}$ on each devices, where $|{\psi}\rangle$ is sampled from a 44-design state ensemble.
4Measure $(\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtains result $|{\phi_{A}}\rangle$.
5Measure $(\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtains result $|{\phi_{B}}\rangle$.
Compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$ and return
$=\int\operatorname{Tr}\left[\left(\frac{I+TU|{\psi}\rangle\!\langle{\psi}|U^{\dagger}}{d+T}\right)\left(\frac{I+TV|{\psi}\rangle\!\langle{\psi}|V^{\dagger}}{d+T}\right)\right]{\rm d}\psi$ Lemma 10
$=\frac{d+2T}{(d+T)^{2}}+\frac{T^{2}}{(d+T)^{2}}\int\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]{\rm d}\psi$ (61)
As we can see, the proof relies on Lemma 10, which requires that the output states remain pure. For a general quantum channel, the output states might be mixed, and the key identities in the lemma no longer hold. Therefore, the coherent algorithm is applicable only when the unknown channels are unitary channels. Lastly, we can prove $\tilde{\chi}$ is an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$,
$\mathbb{E}\tilde{\chi}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\mathbb{E}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}.$ (62)
$X=\frac{(d+1)(d+T)^{2}}{T^{2}d},\quad Y=\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (63)
Then, we have
[…]
It is suffices to consider $\mathbb{E}\tilde{f}^{2}$. With [15, Proof of Lemma 5, Eqs. (165) and (166)], we have
where we define
With Lemma 7, we have
and
$=\frac{1}{d(d+1)(d+2)(d+3)}\sum_{\pi\in\mathcal{S}_{4}}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}P_{\pi}\right]\leq\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}+\mathcal{O}\left(\frac{1}{d}\right)$ (69)
Thus, the variance of $\tilde{\chi}$ is upper bounded by
$\mathbb{V}(\tilde{\chi})\leq\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d}{T^{3}}+\frac{d^{2}}{T^{4}}\right)+\mathcal{O}\left(\frac{d}{T^{2}}+\frac{1}{T}+\frac{1}{d}\right)=\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d^{2}}{T^{4}}\right).$ (70)
Therefore, to achieve $\varepsilon$ additive error, we require
Lemma 10 (Lemmas 13 and 14 in [42]).
For measurement $\mathcal{M}_{s}$ on pure state $|{\psi}\rangle\!\langle{\psi}|^{\otimes s}$, the expectation of measurement result is
Additionally, we have
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|^{\otimes 2}=\frac{2}{(d+s)(d+s+1)}\left[(I+s|{\psi}\rangle\!\langle{\psi}|)^{\otimes 2}-\frac{s(s+1)}{2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]\Pi_{\rm sym}^{(2)}.$ (73)
Appendix C Lower Bound
In this section, we establish lower bounds for completing DSEU with both incoherent and coherent accesses. Here, we consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $n$-qubit subsystems of a Hilbert space $\mathcal{H}\simeq\mathcal{H}_{\rm main}\otimes\mathcal{H}_{\rm aux}$, where $\mathcal{H}_{\rm main}\simeq(\mathbb{C}^{2})^{\otimes n}$ is the ‘main system’ comprising $n$ qubits and $\mathcal{H}_{\rm aux}\simeq(\mathbb{C}^{2})^{\otimes n^{\prime}}$ is an ‘auxiliary system’ of $n^{\prime}$ qubits. It is convenient to define $d=2^{n}$ and $d^{\prime}=2^{n^{\prime}}$. Additionally, as pointed in [37, Remark 4.19], we only need to consider rank-$1$ POVM on two quantum systems.
Our general approach is to reduce the prediction task to a two-hypothesis distinguishing problem. In particular, to find the lower bound of completing DSEU, we consider the following distinguishing problem.
@@ Appendix C Lower Bound
If a learning algorithm can estimate values of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$, then, we can use this learning algorithm to complete the above distinguishing problem. Hence, a lower bound for this distinguishing problem also gives a lower bound for DSEU. To bound the success probability of solving the above distinguishing problem, we use Le Cam’s two-point method as follows.
@@
In the remainder of this section, we analyze TVD between $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$, respectively, thereby establishing the corresponding lower bounds.
C.1 Incoherent Access
Here we will prove the lower bound for completing DSEU with incoherent access. First, we introduce the tree representation of channel learning with incoherent access, which is a powerful tool in proving the lower bound of learning task [37, 38, 44, 45, 46].
C.1.1 Proof of Theorem 3
@@
Consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $n$-qubit subsystems of a Hilbert space $\mathcal{H}$. A learning algorithm with incoherent access can be represented as a rooted tree $\mathcal{T}$ of depth $T$ such that each node encodes all measurement outcomes the algorithm has received thus far. The tree has the following properties:
@@
  • At each non-leaf node $u$, we prepare a state $|{\phi_{u,1}}\rangle\otimes|{\phi_{u,2}}\rangle$ on $\mathcal{H}^{\otimes 2}$, apply channels $\mathcal{U}$ and $\mathcal{V}$ onto two $n$-qubit subsystems, and measure a rank-$1$ POVM $\{w_{v}^{u}d^{2}d^{\prime 2}|{\psi^{u}_{v,1}}\rangle\!\langle{\psi^{u}_{v,1}}|\otimes|{\psi^{u}_{v,2}}\rangle\!\langle{\psi^{u}_{v,2}}|\}_{v}$ (which can depend on $u$) on the entire system to obtain a classical outcome $v$. Each child node $v$ of the node $u$ corresponds to a particular POVM outcome $v$ and is connected by the edge $e_{u,v}$. We refer to the set of child node of node $u$ a ${\rm child}(u)$.
@@
  • Each root-to-leaf path is of length $T$. For a leaf of corresponding to node $\ell$, $p^{U,V}(\ell)$ is the probability that the classical memory is in state $\ell$ after the learning procedure.
Based on the above tree representation and distinguishing problem, we are ready to prove the lower bound for completing DSEU with incoherent access, i.e., Theorem 3. As shown in Lemma 11, our goal is to obtain an upper bound for $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$. With triangle inequality, we have the following upper bound for each leaf $\ell$,
@@
where $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$ is the completely depolarizing channel and $p^{\mathcal{D}}(\ell)$ is the probability that both two quantum channels are $\mathcal{D}$. With Lemma 13, we can find the upper bounds for two terms in Eq. (77) and have
@@
C.1.2 Technical Lemmas

### v2

@@ A.1 Notation
A.2 Haar Measure and Permutation Operators
@@
Let $\mathcal{H}$ be a Hilbert space with dimension $d$, $\mathcal{S}_{k}$ be the symmetric group, and $A\in\mathcal{B}(\mathcal{H}^{\otimes k})$. Then, we have
@@
For quantum states in $\mathcal{H}$, we define the Haar measure on quantum states as [56]
$\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}|{\phi}\rangle\!\langle{\phi}|^{\otimes k}U^{\dagger\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}},\quad|{\phi}\rangle\!\langle{\phi}|\in\mathcal{D}(\mathcal{H}),$ (22)
@@
$\mathbb{E}_{|{\psi}\rangle\sim\mathcal{A}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}=\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}}.$ (24)
@@
Based on the properties of state designs, we have the following results for the symmetric collective measurement $\mathcal{M}_{s}$.
Lemma 9 (Lemmas 13 and 14 in [52]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$. For the measurement $\mathcal{M}_{s}$ on the pure state $|{\psi}\rangle\!\langle{\psi}|^{\otimes s}$, the expectation of the measurement result is
Additionally, we have
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|^{\otimes 2}=\frac{2}{(d+s)(d+s+1)}\left[(I+s|{\psi}\rangle\!\langle{\psi}|)^{\otimes 2}-\frac{s(s+1)}{2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]\Pi_{\rm sym}^{(d,2)}.$ (27)
A.3 Approximate Unitary Designs
@@
Appendix B Lower Bound
In this section, we establish lower bounds for completing DSEC with both incoherent and coherent accesses. Here, we consider two unitary channels $\mathcal{U}=U(\cdot)U^{\dagger}$ and $\mathcal{V}=V(\cdot)V^{\dagger}$ acting on $d$-dimensional subsystems of a Hilbert space $\mathcal{H}\simeq\mathcal{H}_{\rm main}\otimes\mathcal{H}_{\rm aux}$, where $\mathcal{H}_{\rm main}$ is the main system with dimension $d$ and $\mathcal{H}_{\rm aux}$ is an auxiliary system with dimension $d^{\prime}$. Additionally, as shown in [44, Remark 4.19], we only need to consider rank-$1$ POVMs on two quantum systems.
Our approach is to reduce the estimation task to a two-hypothesis distinguishing problem. In particular, to derive a lower bound for completing DSEC, we consider the following distinguishing problem, which is a restatement of Problem 1 in the main text:
@@ Appendix B Lower Bound
Define the Choi operators of $U$ and $V$ as $J_{U}$ and $J_{V}$, respectively. If an estimation algorithm can accurately estimate $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$, then it can be leveraged to solve the distinguishing problem above. Consequently, any lower bound for this distinguishing problem immediately yields a lower bound for DSEC. To bound the success probability of solving the distinguishing problem, we invoke Le Cam’s two-point method as follows.
@@
In the remainder of this section, we analyze the TVD between $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$, respectively, thereby establishing the corresponding lower bounds.
B.1 Incoherent Access
Here, we prove the lower bound for completing DSEC with incoherent access. First, we introduce the tree representation of channel learning with incoherent access, which is a powerful tool in proving lower bounds for learning tasks [44, 41, 47, 48, 49].
B.1.1 Proof of Theorem 1
@@
Consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $\mathcal{H}_{\rm main}$. A learning algorithm with incoherent access can be represented as a rooted tree $\mathcal{T}$ of depth $T$ such that each node encodes all measurement outcomes the algorithm has received thus far. The tree has the following properties:
@@
  • At each non-leaf node $u$, we prepare a state $|{\phi_{u,1}}\rangle\otimes|{\phi_{u,2}}\rangle$ on $\mathcal{H}^{\otimes 2}$, apply channels $\mathcal{U}$ and $\mathcal{V}$ onto two $d$-dimensional subsystems, and measure a rank-$1$ POVM $\{w_{v}^{u}d^{2}d^{\prime 2}|{\psi^{u}_{v,1}}\rangle\!\langle{\psi^{u}_{v,1}}|\otimes|{\psi^{u}_{v,2}}\rangle\!\langle{\psi^{u}_{v,2}}|\}_{v}$ (which can depend on $u$) on the entire system to obtain a classical outcome $v$. Each child node $v$ of the node $u$ corresponds to a particular POVM outcome $v$ and is connected by the edge $e_{u,v}$. We denote the set of child nodes of node $u$ as ${\rm child}(u)$.
@@
  • Each root-to-leaf path is of length $T$. For a leaf corresponding to node $\ell$, $p^{U,V}(\ell)$ is the probability that the classical memory is in state $\ell$ after the learning procedure.
Based on the above tree representation and distinguishing problem, we are ready to prove the lower bound for completing DSEC with incoherent access, i.e., Theorem 1. As shown in Lemma 11, our goal is to obtain an upper bound for $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$. By the triangle inequality, we have the following upper bound for each leaf $\ell$,
@@
where $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$ is the completely depolarizing channel and $p^{\mathcal{D}}(\ell)$ is the probability that both two quantum channels are $\mathcal{D}$. Using Lemma 13, we can find upper bounds for the two terms in Eq. (35) and obtain
@@
which implies that we require $T=\Omega(\sqrt{d})$. Therefore, we complete the proof of Theorem 1.
B.1.2 Technical Lemmas